Two people start driving toward each other on the same highway at the same time from 420 miles apart. one car travels 15 mph faster then the other. what is the rate of each car if they meet in 4 hrs.


answer: the first car travels 60 mph; the second travels at 45 mph i got as far as d1 + d2= 420 ~~~~~> d2=420 - d1
r2=r1 + 15 need a full break down of answer

Answer by Mathtut(3670)

d=r*t
:
d=(r+15)4...eq 1---->d=4r+60.......fast car
420-d=r(4)..eq 2---->d=-4r+420.....slow car
:
4r+60=-4r+420
:
8r=360
:
r=45 slow car
:
r+15=60...fast car
: